If an object in space has a radius of 5x106m and a mass of 8x1023kg what is the acceleration due to gravity on the surface of the object?
Answer #1

acceleration on the earth is given by g = GM/R^2; M is the mass and R is radius

R = 5*10^6 m

M = 8*1023 kg

using this we get

g = 2.136 m/sec2